# Breather-to-soliton transitions and nonlinear wave interactions for the higher-order modified Gerdjikov–Ivanov equation

### v1

@@ 2.1 $\frac{V_{1,H}}{H_{1,R}}\neq\frac{V_{1,T}}{H_{1,I}}$
In this case, the first-order solution become the localized wave solution with breathing behavior. Under the condition $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0$, when $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)>$ 0, the general breather solution is obtained as shown in Fig.(1(a)). The Akhmediev breather solution which is localized in time and periodic in space is presented in Fig.(1(b)) when the condition $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)<$ 0 holds.
@@
Figure 1: The first-order breather solution. (a) The general breather solution with $\lambda_{1}=\frac{1}{3}+\frac{1}{2}\mathrm{i},a=\frac{3}{10},c=1,\delta=\frac{1}{10},c_{1}=c_{2}=2,b=-\frac{11}{50}$; (b) The Akhmediev breather solution with $\lambda_{1}=\frac{4}{5}+\frac{\sqrt{14}}{10}\mathrm{i},a=\frac{3}{5},c=1,\delta=-\frac{1}{5},c_{1}=c_{2}=1,b=-\frac{751}{1000}$.
2.2 $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$
@@
$-2c^{2}\delta+2\delta-\frac{1}{2}c^{2}-2\beta_{1}^{2}+2\alpha_{1}^{2}+a=0.$ (4)
Given that the condition (4) is satisfied, we can obtain the multi-peak soliton as is shown in Fig.2 by choosing appropriate parameters in Eq.(3) when $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}\neq 0$. We can find that Fig.(2(a))-Fig.(2(d)) are the oscillation W-shaped soliton, while Fig.(2(e))-Fig.(2(h)) are the oscillation M-shaped soliton. The transformation between the two kinds of solitons can be finished by changing the sign of the imaginary part $\beta_{1}$ of the eigenvalue. Futhermore, it’s observed that when the real part $\alpha_{1}>0$ of the eigenvalue decreases, the number of wave peaks for the oscillation W-shaped and M-shaped soliton will increases.
@@
Figure 2: The multi-peak soliton solution with $a=\frac{2}{5},c=1,\delta=\frac{7}{10},c_{1}=c_{2}=2,b=-\frac{5109}{1000}$. (a) The oscillation W-shaped soliton with $\lambda_{1}=\frac{2}{5}-\frac{\sqrt{11}}{10}\mathrm{i}$; (b) The cross-sectional view of (a) at $t=0$; (c) The oscillation W-shaped soliton with $\lambda_{1}=\frac{3}{10}-\frac{\sqrt{1}}{5}\mathrm{i}$; (d) The cross-sectional view of (c) at $t=0$; (e) The oscillation M-shaped soliton with $\lambda_{1}=\frac{2}{5}+\frac{\sqrt{11}}{10}\mathrm{i}$; (f) The cross-sectional view of (e) at $t=0$; (g) The oscillation M-shaped soliton with $\lambda_{1}=\frac{3}{10}+\frac{\sqrt{1}}{5}\mathrm{i}$; (h) The cross-sectional view of (g) at $t=0$.
@@
Based on the condition (4), unlike the multi-peak soliton, we make another condition $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0$ holds. In this case, the periodic wave solution and anti-dark soliton solution are obtained. We can derive two different types of periodic wave solutions as is shown in Fig.3 by choosing appropriate parameters when $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)<$ 0. It is worth noting that the period of the periodic wave is $\frac{2\pi}{H_{I}}$, hence when the period extend to infinity, namely $H_{I}\to 0$, the periodic wave solution will be converted to the W-shape soliton solution as is presented in Fig.(4(a))-(4(b)). Subsequently, when $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)>$ 0, the first-order solution can be degenerate into the anti-dark soliton solution which is a bell-shaped localized wave on the plane wave background in Fig.(4(c))-(4(d)).
@@
Figure 3: The periodic wave solution with $a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$. (a) $\lambda=\frac{1}{2}+\frac{\sqrt{15}}{10}\mathrm{i}$; (b) The cross-sectional view of (a) at $t=0$; (c) $\lambda=\frac{7}{20}+\frac{3}{20}\mathrm{i}$; (d) The cross-sectional view of (d) at $t=0$.
@@
Figure 4: (a)The W-shaped soliton degenerated from the periodic wave Fig.(3(a)) with $\lambda_{1}=\frac{\sqrt{35}}{10}-\frac{1}{2}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{50},c_{1}=c_{2}=1,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at $t=0$; (c) The anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{59}}{10}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (d) The cross-sectional view of (c) at $t=0$.
@@ 3 Nonlinear wave interactions
In this section, the breather-to-soliton state transition of the second-order solution is discussed. In other words, we will explore the nonlinear converted wave interactions. We choose $n=2$, $H_{2}=H_{2,R}+\mathrm{i}H_{2,I}$ and $\lambda_{2}=\alpha_{2}+\mathrm{i}\beta_{2}$, then the second-order solution contains an additional hyperbolic function part $\cosh(V_{2,H}t+H_{2,R}x),~{}\sinh(V_{2,H}t+H_{2,R}x)$ and a trigonometric function part $\cos(V_{2,T}t+H_{2,I}x),~{}\sin(V_{2,T}t+H_{2,I}x)$ compared to the first-order solution. Hence, we need to consider two state transition conditions, namely $\frac{V_{j,H}}{H_{j,R}}=\frac{V_{j,T}}{H_{j,I}},j=1,2$ or $\frac{V_{j,H}}{H_{j,R}}\neq\frac{V_{j,T}}{H_{j,I}},j=1,2$.
@@
Firstly, under the conditions $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$ and $\frac{V_{2,H}}{H_{2,R}}\neq\frac{V_{2,T}}{H_{2,I}}$, the interactions between the breather and other nonlinear converted waves can be obtained. Fig.5 exhibits the interaction between the breather and the anti-dark soliton when additional conditions $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0$ and $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)>$ 0 are satisfied. When $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)<$ 0 holds, we can obtain the interaction between the breather and the periodic wave as is shown in Fig.6. Moreover, when the condition $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0$ is satisfied, the second-order solution can be degenerated into the interaction between the breather and multi-peak soliton and we can see Fig.7.
@@
Figure 5: The interaction between the breather and the anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{59}}{10}\mathrm{i},\lambda_{2}=1+\frac{1}{5}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
@@
Figure 6: The interaction between the breather and the periodic wave with $\lambda_{1}=\frac{7}{20}-\frac{3}{20}\mathrm{i},\lambda_{2}=1-\frac{1}{5}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
@@
Figure 7: The interaction between the breather and the multi-peak soliton with $\lambda_{1}=\frac{2}{5}-\frac{\sqrt{11}}{10}\mathrm{i},\lambda_{2}=-\frac{1}{2}+\mathrm{i},a=\frac{2}{5},c=1,\delta=\frac{7}{10},c_{1}=c_{2}=2,b=-\frac{5109}{1000}$; (b) The density plot for (a).
@@
Subsequently, based on the conditions $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$ and $\frac{V_{2,H}}{H_{2,R}}=\frac{V_{2,T}}{H_{2,I}}$, when the conditions $\alpha_{j}^{2}-\beta_{j}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0,j=1,2$ hold and $24\beta_{j}^{2}\alpha_{j}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right),j=1,$ 2 show the same signs, we can obtain the interaction about the W-shaped soliton and the anti-dark soliton by adjusting the parameters. In Fig.8, the interaction between the W-shaped soliton and the W-shaped soliton is obtained. The interaction between the W-shaped soliton and the anti-dark soliton is presented in Fig.9. Meanwhile, we can also derive the interaction between the anti-dark soliton and the anti-dark soliton in Fig.LABEL:ffg3.
@@
Figure 8: The interaction between the W-shaped soliton and the W-shaped soliton with $\lambda_{1}=\frac{7}{10}-\frac{\sqrt{39}}{10}\mathrm{i},\lambda_{2}=-\frac{4}{5}+\frac{3\sqrt{6}}{10}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at $t=0$.
@@
Figure 9: The interaction between the anti-dark soliton and the W-shaped soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i},\lambda_{2}=-\frac{4}{5}+\frac{3\sqrt{6}}{10}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at $t=0$.
@@
Figure 10: The interaction between the anti-dark soliton and the anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i},\lambda_{2}=\frac{4}{5}-\frac{3\sqrt{6}}{10}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at $t=0$.
@@ 4 Double-pole breather-to-soliton transitions
where $\phi_{1},\psi_{1}$ are the same as Eq.(3).
Similarly, the state transition condition of the double-pole solution is $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$. Therefore, based on the condition, we will obtain the double-pole anti-dark soliton, double-pole multi-peak soliton and double-pole periodic wave solution. When the additional condition $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}\neq 0$ is satisfied, the double-pole solution can be reduced into the double-pole multi-peak soliton solution. We can see that Fig.11 presents the interaction between two series of the double-pole multi-peak soliton. Their wave peaks achieve the maximal amplitude at $(0,0)$. If the condition $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0$ holds, when $4\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)<0(>0)$, we can obtain the double-pole periodic wave solution and the double-pole anti-dark soliton solution as is shown in Fig.12 and Fig.13, respectively. The double-pole periodic wave solution possesses a periodic chained high-amplitude region which exhibits the double-pole phenomena.
Figure 11: (a) The double-pole multi-peak soliton solution with $\lambda_{1}=\frac{1}{2}-\frac{\sqrt{5}}{5}\mathrm{i},a=\frac{2}{5},c=1,\delta=\frac{7}{10},c_{1}=c_{2}=2,b=-\frac{5109}{1000}$; (b) The density plot for (a).
Figure 12: (a) The double-pole periodic wave solution with $\lambda_{1}=\frac{1}{2}-\frac{\sqrt{15}}{5}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
Figure 13: (a) The double-pole anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
We will apply the asymptotic analysis method described in [28] to explore the asymptotic behavior of the double-pole anti-dark soliton. For convenience, assume $c=1,c_{1}=c_{2}=2,\lambda_{1}=\alpha_{1}+\mathrm{i}\beta_{1}$ that under the corresponding state transition conditions, we can obtain $a=\alpha_{1}^{2}-\beta_{1}^{2}+\frac{4\delta+1}{2},\delta=-\frac{3\alpha_{1}^{2}}{2}+\frac{3\beta_{1}^{2}}{2}$ and $14\beta_{1}^{2}>$ 1. Hence, the double-pole anti-dark soliton solution is
@@
Subsequently, we substitute $x=\frac{\left(6\beta_{1}^{4}+4(5\alpha_{1}^{2}-1)\beta_{1}^{2}+6\alpha_{1}^{4}+4\alpha_{1}^{2}+1\right)t-D_{2}}{4}$ into Eq.(6) so that the double-pole solution is only related to $D_{2}$ and $t$. Numerous articles about the asymptotic analysis method have illustrated that the characteristic line of the double-pole solution is curved[29, 30, 31]. Therefore, according to the relation $e^{D_{2}}\geq D_{2}(D_{2}\to+\infty)$ and $e^{D_{2}}\leq D_{2}(D_{2}\to-\infty)$, there must exist an asymptotic balance relation between $e^{D_{2}}$ and $t$. The explicit expressions of the asymptotic solitons are exhibited as follows.
$\bullet$  When $e^{D_{2}}\sim t$, $D_{2}\to+\infty$,

### v2

@@ 2.1 The case $\frac{V_{1,H}}{H_{1,R}}\neq\frac{V_{1,T}}{H_{1,I}}$
the general breather solution is obtained as shown in Fig.(1(a)). The Akhmediev breather solution which is localized in time and periodic in space is presented in Fig.(1(b)) when adding the condition
$4\beta_{1}^{2}\alpha_{1}^{2}+\frac{d^{4}}{4}+d^{2}\left(2d^{2}\delta-a\right)<0.$ (7)
@@
Figure 1: The first-order breather solution. (a) The spatio-temporal periodic breather solution with $\lambda_{1}=\frac{1}{3}+\frac{1}{2}\mathrm{i},a=\frac{3}{10},d=1,\delta=\frac{1}{10},c_{1}=c_{2}=2,b=-0.22$; (b) The Akhmediev breather solution with $\lambda_{1}=\frac{4}{5}+\frac{\sqrt{14}}{10}\mathrm{i},a=\frac{3}{5},d=1,\delta=-\frac{1}{5},c_{1}=c_{2}=1,b=-0.751$.
2.2 The case $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$
@@
$-2d^{2}\delta+2\delta-\frac{1}{2}d^{2}-2\beta_{1}^{2}+2\alpha_{1}^{2}+a=0.$ (8)
Given that the condition (8) is satisfied, we can obtain the multi-peak soliton as is shown in Fig.2 by choosing appropriate parameters in Eq.(4) when
$\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)d^{2}-2a}{2}\neq 0.$ (9)
We can find that Fig.(2(a))-Fig.(2(d)) are the oscillation W-shaped soliton, while Fig.(2(e))-Fig.(2(h)) are the oscillation M-shaped soliton. Reversing the sign of the imaginary part $\beta_{1}$ of the eigenvalue leads to the transition between these two solitons. Unlike the general solitons, the energy of multi-peak solitons is distributed among multiple separated localized peaks, and the stability between these peaks is maintained through a balance between nonlinearity and dispersion. Furthermore, it is observed that when $\alpha_{1}>0$ decreases, the number of wave peaks for the oscillation W-shaped and M-shaped soliton will increase.
@@
Figure 2: The multi-peak soliton solution with $a=\frac{2}{5},d=1,\delta=\frac{7}{10},c_{1}=c_{2}=2,b=-5.109$. (a) The oscillation W-shaped soliton with $\lambda_{1}=\frac{2}{5}-\frac{\sqrt{11}}{10}\mathrm{i}$; (b) The cross-sectional view of (a) at $t=0$; (c) The oscillation W-shaped soliton with $\lambda_{1}=\frac{3}{10}-\frac{1}{5}\mathrm{i}$; (d) The cross-sectional view of (c) at $t=0$; (e) The oscillation M-shaped soliton with $\lambda_{1}=\frac{2}{5}+\frac{\sqrt{11}}{10}\mathrm{i}$; (f) The cross-sectional view of (e) at $t=0$; (g) The oscillation M-shaped soliton with $\lambda_{1}=\frac{3}{10}+\frac{1}{5}\mathrm{i}$; (h) The cross-sectional view of (g) at $t=0$.
@@
Based on the condition (8), unlike the multi-peak soliton, we make another condition (5) holds. In this case, the periodic wave solution and anti-dark soliton solution are obtained. We can derive two different types of periodic wave solutions as is shown in Fig.3 by choosing appropriate parameters when (7). It is worth noting that the period of the periodic wave is $\frac{2\pi}{H_{I}}$, hence when the period extends to infinity, namely $H_{I}\to 0$, the periodic wave solution will be converted to the W-shaped soliton solution as is presented in Fig.(4(a))-(4(b)). The W-shaped soliton solution is rational without the hyperbolic function and the trigonometric function, we show its explicit expression as follows,
$q_{rs}=\frac{-213000[-\frac{639}{1000}t^{2}+(\mathrm{i}+\frac{6}{5}x)t-\frac{200\mathrm{i}}{213}x-\frac{40}{71}x^{2}+\frac{200}{71}]e^{-\frac{\mathrm{i}}{160}(t-48x)}}{-136107t^{2}+(213000\mathrm{i}+255600x)t-200000\mathrm{i}x-120000x^{2}-200000}.$
Subsequently, when the codition (6) holds, the first-order solution can be degenerate into the anti-dark soliton solution which is a bell-shaped localized wave on the plane wave background in Fig.(4(c))-(4(d)).
@@
Figure 3: The periodic wave solution with $a=\frac{3}{10},d=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$. (a) $\lambda=\frac{1}{2}+\frac{\sqrt{15}}{10}\mathrm{i}$; (b) The cross-sectional view of (a) at $t=0$; (c) $\lambda=\frac{7}{20}+\frac{3}{20}\mathrm{i}$; (d) The cross-sectional view of (d) at $t=0$.
@@
Figure 4: (a)The W-shaped soliton degenerated from the periodic wave Fig.(3(a)) with $\lambda_{1}=\frac{\sqrt{35}}{10}-\frac{1}{2}\mathrm{i},a=\frac{3}{10},d=1,\delta=-\frac{3}{50},c_{1}=c_{2}=1,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at $t=0$; (c) The anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{59}}{10}\mathrm{i},a=\frac{3}{10},d=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (d) The cross-sectional view of (c) at $t=0$.
@@ 3 Nonlinear wave interactions
In this section, the breather-to-soliton state transition of the second-order solution is discussed. In other words, we will explore the nonlinear converted wave interactions. When $n=2$, we have $H_{2}=H_{2,R}+\mathrm{i}H_{2,I}$ which corresponds to the spectral parameter $\lambda_{2}=\alpha_{2}+\mathrm{i}\beta_{2}$. Then the second-order solution contains an additional hyperbolic function part $\cosh(V_{2,H}t+H_{2,R}x),~\sinh(V_{2,H}t+H_{2,R}x)$ and a trigonometric function part $\cos(V_{2,T}t+H_{2,I}x),~\sin(V_{2,T}t+H_{2,I}x)$ compared to the first-order solution. Thus, due to the existence of two eigenvalue parameters, in addition to considering the condition of the case 2.1 or the case 2.2, it is necessary to take into account another condition, namely $\frac{V_{2,H}}{H_{2,R}}=\frac{V_{2,T}}{H_{2,I}}$ or $\frac{V_{2,H}}{H_{2,R}}\neq\frac{V_{2,T}}{H_{2,I}}$.
@@
Firstly, under the conditions $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$ and $\frac{V_{2,H}}{H_{2,R}}\neq\frac{V_{2,T}}{H_{2,I}}$, the interactions between the breather and other nonlinear converted waves can be obtained. Fig.5 exhibits the interaction between the breather and the anti-dark soliton when additional conditions (5) and (6) are satisfied. When the condition (7) holds, we can obtain the interaction between the breather and the periodic wave as is shown in Fig.6. Moreover, when the condition (5) is satisfied, the second-order solution can be degenerated into the interaction between the breather and multi-peak soliton and we can see Fig.7.
@@
Figure 5: The interaction between the breather and the anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{59}}{10}\mathrm{i},\lambda_{2}=1+\frac{1}{5}\mathrm{i},a=\frac{3}{10},d=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
@@
Figure 6: The interaction between the breather and the periodic wave with $\lambda_{1}=\frac{7}{20}-\frac{3}{20}\mathrm{i},\lambda_{2}=1-\frac{1}{5}\mathrm{i},a=\frac{3}{10},d=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
@@
Figure 7: The interaction between the breather and the multi-peak soliton with $\lambda_{1}=\frac{2}{5}-\frac{\sqrt{11}}{10}\mathrm{i},\lambda_{2}=-\frac{1}{2}+\mathrm{i},a=\frac{2}{5},d=1,\delta=\frac{7}{10},c_{1}=c_{2}=2,b=-5.109$; (b) The density plot for (a).
@@
Subsequently, based on the conditions $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$ and $\frac{V_{2,H}}{H_{2,R}}=\frac{V_{2,T}}{H_{2,I}}$, when the conditions $\alpha_{j}^{2}-\beta_{j}^{2}+\frac{(4\delta+1)d^{2}-2a}{2}=0,j=1,2$ hold and $4\beta_{j}^{2}\alpha_{j}^{2}+\frac{d^{4}}{4}+d^{2}\left(2d^{2}\delta-a\right),j=1,2$ show the same signs, we can obtain the interaction about the W-shaped soliton and the anti-dark soliton by adjusting the parameters. In Fig.8, the interaction between the W-shaped soliton and the W-shaped soliton is obtained. The interaction between the W-shaped soliton and the anti-dark soliton is presented in Fig.9. Meanwhile, we can also derive the interaction between the anti-dark soliton and the anti-dark soliton in Fig.10.
@@
Figure 8: The interaction between the W-shaped soliton and the W-shaped soliton with $\lambda_{1}=\frac{7}{10}-\frac{\sqrt{39}}{10}\mathrm{i},\lambda_{2}=-\frac{4}{5}+\frac{3\sqrt{6}}{10}\mathrm{i},a=\frac{3}{10},d=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at t=20,0,20t=-20,0,20.
@@
Figure 9: The interaction between the anti-dark soliton and the W-shaped soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i},\lambda_{2}=-\frac{4}{5}+\frac{3\sqrt{6}}{10}\mathrm{i},a=\frac{3}{10},d=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at t=20,0,20t=-20,0,20.
@@
Figure 10: The interaction between the anti-dark soliton and the anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i},\lambda_{2}=\frac{4}{5}-\frac{3\sqrt{6}}{10}\mathrm{i},a=\frac{3}{10},d=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at t=20,0,20t=-20,0,20.
@@ 4 Double-pole breather-to-soliton transitions
where $\epsilon$ is a small parameter and $\phi_{1},\psi_{1}$ are the same as in Eq.(4).
Similarly, the state transition condition of the double-pole solution is $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$. Therefore, based on the condition, we will obtain the double-pole anti-dark soliton, double-pole multi-peak soliton and double-pole periodic wave solution. When the additional condition (9) is satisfied, the double-pole solution can be reduced into the double-pole multi-peak soliton solution. We can see that Fig.11 presents the interaction between two series of the double-pole multi-peak soliton. Their wave peaks achieve the maximal amplitude at $(0,0)$. If the condition (5) holds, when (7) and (6) are satisfied, we can obtain the double-pole periodic wave solution and the double-pole anti-dark soliton solution as is shown in Fig.12 and Fig.13, respectively. The double-pole periodic wave solution possesses a periodic chained higher-amplitude region which exhibits the double-pole phenomena.
Figure 11: (a) The double-pole multi-peak soliton solution with $\lambda_{1}=\frac{1}{2}-\frac{\sqrt{5}}{5}\mathrm{i},a=\frac{2}{5},d=1,\delta=\frac{7}{10},c_{1}=c_{2}=2,b=-5.109$; (b) The density plot for (a).
Figure 12: (a) The double-pole periodic wave solution with $\lambda_{1}=\frac{1}{2}-\frac{\sqrt{15}}{5}\mathrm{i},a=\frac{3}{10},d=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
Figure 13: (a) The double-pole anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i},a=\frac{3}{10},d=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
We will apply the asymptotic analysis method described in [39] to explore the asymptotic behavior of the double-pole anti-dark soliton. For convenience, assuming $d=1,c_{1}=c_{2}=2,\lambda_{1}=\alpha_{1}+\mathrm{i}\beta_{1}$ under the corresponding state transition conditions, we can obtain $a=\alpha_{1}^{2}-\beta_{1}^{2}+\frac{4\delta+1}{2},\delta=-\frac{3\alpha_{1}^{2}}{2}+\frac{3\beta_{1}^{2}}{2}$ and $4\beta_{1}^{2}>1$. Hence, the double-pole anti-dark soliton solution is
@@
Subsequently, we substitute $x=\frac{\left(6\beta_{1}^{4}+4(5\alpha_{1}^{2}-1)\beta_{1}^{2}+6\alpha_{1}^{4}+4\alpha_{1}^{2}+1\right)t-D_{2}}{4}$ into Eq.(11) so that the expression of the double-pole solution is only related to $D_{2}$ and $t$. Numerous articles about the asymptotic analysis method have illustrated that the characteristic line of the double-pole solution is curved [39, 40, 41]. Therefore, according to the asymptotic analysis method, if we want to derive the asymptotic soliton, an asymptotic balance relation between $e^{\pm D_{2}}$ and $t$ need to be satisfied [42]. The explicit expressions of the asymptotic solitons in general form are exhibited as follows,
$\bullet$  When $e^{D_{2}}\sim t$,